# ImagineNav++: Prompting Vision-Language Models as Embodied Navigator through Scene Imagination

### v1

@@ IV-D Numerical Evaluations on InsINav Task
Table II summarizes the performance of our ImagineNav++ and baseline methods on the InsINav task. While UniGoal [yin2025unigoaluniversalzeroshotgoaloriented] achieves the highest SR of 60.2%, it requires a graph-based planner that introduces significant computational overhead from graph construction, graph update, and graph matching. In comparison, our ImagineNav++ operates under a strict mapless setting and still attains a competitive SR of 52.4%. More notably, ImagineNav++ establishes a new state-of-the-art in path efficiency and achieves a superior SPL of 32.8%, significantly outperforming all other methods, including those that leverage map information (Mod-IIN [krantz2023navigatingobjectsspecifiedimages]: 23.3%, Goat [chang2024goatthing]: 24.1%, UniGoal [yin2025unigoaluniversalzeroshotgoaloriented]: 23.7%). This advantage stems from our proactive planning paradigm, wherein the agent utilizes an imagination mechanism to synthesize and evaluate potential gains across multiple future directions, and subsequently integrates historical observations to select the optimal exploratory path, thereby minimizing haphazard exploration and redundant detours. This demonstrates that our ImagineNav++ achieves a superior navigation efficiency, excelling in both final goal attainment and the optimality of the paths taken. Among methods adhering to the same open-vocabulary, mapless, and universal constraints, our ImagineNav++ delivers markedly better performance than PSL [sun2024prioritizedsemanticlearningzeroshot], more than doubling both SR (52.4% vs. 23.0%) and SPL (32.8% vs. 11.4%). These results highlight the capability of our ImagineNav++ to enable efficient and robust navigation relying solely on visual observations, without requiring explicit map construction. The efficacy of our framework on instance‑level objectives further validates the strong adaptability of our ImagineNav++ across diverse goal‑oriented visual navigation tasks.
@@ IV-F1 Choice of Imagination Horizon
Figure 5: Where2Imagine: the impact of different sampling interval $\mathbf{T}$ on ObjectNav performance on HM3D dataset.
@@
We evaluate the impact of different visual backbones on both waypoint prediction and final navigation performance. All backbones were adapted for our task: ResNet-18 and ViT were modified at their final layers and trained end-to-end, whereas DINOv2 and MAE were kept frozen, with their features fed into a separate trainable five-layer MLP. All experiments used real RGB observations without resorting to the NVS model and were tested for 100 epoches. As shown in Table IV, ResNet-18 achieves the best performance in both waypoint prediction and ObjectNav, despite being the most lightweight architecture. Furthermore, the significant performance gap across different backbones underscores the importance of our Where2Imagine module, indicating the usefulness of learning from human demonstrations. We also note a slight discrepancy between the waypoint prediction loss and the navigation success rate for DINOv2 and MAE. This is likely attributable to the domain shift between the MP3D dataset (used for waypoint evaluation) and the HM3D environment (used for navigation), which introduces evaluation variability.
@@
Figure 7 presents a comprehensive qualitative evaluation of our proposed Where2Imagine module on the HM3D dataset. This visualization reveals several key insights. First, the waypoints generated by our Where2Imagine module is concentrated in navigable regions, yielding observation views that are semantically richer and capture essential structural elements such as furniture, doors, and hallways. This stands in clear contrast to the uniform sampling strategy (with a fixed radius of 2.0m and angular interval of $60^{\circ}$), which disregards scene geometry and frequently fails to capture critical semantic information. This observation confirms the capability of our Where2Imagine module for semantic-aware planning, aligning its exploratory behavior with perceptually salient structural cues. Another key finding emerges from comparing the ground-truth and synthesized view observations. Specifically, although the NVS model is prone to introducing artifacts and inaccuracies, notably in the synthesis of fine details, they generally achieve fundamental consistency with the true scene observation in both semantics and overall structure across diverse HM3D environments. This high-level consistency enables effective spatial reasoning in LLMs, which explains the performance gains observed in navigation tasks.
@@ IV-H Analysis of VLM planner
We conducted a comparative evaluation of the effects of different VLMs on navigation performance, as detailed in Table VI. The experiments used real RGB without NVS model. The results demonstrate that GPT-4o-mini [openai2024gpt4omini] and GPT-4-Turbo [openai2023gpt4turbo] exhibit negligible differences in success rate and SPL metrics, while substantially outperforming LLaVA [10655294], demonstrating that advanced reasoning capabilities are crucial for navigation tasks. Furthermore, our results indicate that within the same model family, strategically selecting more cost-effective variants (such as GPT-4o-mini vs. GPT-4-Turbo) achieves comparable navigation performance while dramatically reducing operational costs by approximately 98%. This optimization enables more efficient resource allocation and faster inference without sacrificing navigation performance, providing practical guidance for real-world deployment scenarios where computational efficiency and operational costs are significant concerns.
@@
GPT-4-Turbo [openai2023gpt4turbo] \$20.00 66.0 36.3
GPT-4o-mini [openai2024gpt4omini] \$0.375 70.0 36.9
@@ IV-J Analysis of Successful and Failed Trajectories
As shown in Figure 10, our method enables efficient path planning and navigation toward diverse targets. In particular, the top-middle subfigure illustrates a scenario where the agent must traverse multiple rooms. Although such complex environments increase the risk of disorientation, our ImagineNav++ successfully reaches the goal after exploration, demonstrating its robustness in long-horizon, multi-room settings. Furthermore, as depicted in the upper-right subfigure, the memory module helps detect local search loops from historical observations and autonomously adjust the exploration direction. We also present some failure examples at the bottom of Figure 10. We identified three key factors contributing to these navigation failures. First, the synthesized image from the NVS does not align with the real observation, such as creating objects that are not present in the real scene as shown in the bottom left of Figure 10, which causes the VLM to make incorrect inference. Second, some object instances are neglected for marking by the simulator, and therefore a successfully trajectory is wrongly considered as a failure (a.k.a. false failure) as shown in the middle of Figure 10. Moreover, the visual ambiguity of the target leads directly to recognition failure, causing the agent to terminate the navigation task unsuccessfully due to the lack of a key basis for the stopping decision, as shown in the bottom right of Figure 10.

### v2

@@ IV-D Numerical Evaluations on InsINav Task
Table II summarizes the performance of our ImagineNav++ and baseline methods on the InsINav task. While UniGoal [84] achieves the highest SR of 60.2%, it requires a graph-based planner that introduces significant computational overhead from graph construction, graph update, and graph matching. In comparison, our ImagineNav++ operates under a strict mapless setting and still attains a competitive SR of 52.4%. More notably, ImagineNav++ establishes a new state-of-the-art in path efficiency and achieves a superior SPL of 32.8%, significantly outperforming all other methods, including those that leverage map information (Mod-IIN [34]: 23.3%, Goat [9]: 24.1%, UniGoal [84]: 23.7%). This advantage stems from our proactive planning paradigm, wherein the agent utilizes an imagination mechanism to synthesize and evaluate potential gains across multiple future directions, and subsequently integrates historical observations to select the optimal exploratory path, thereby minimizing haphazard exploration and redundant detours. This demonstrates that our ImagineNav++ achieves a superior navigation efficiency, excelling in both final goal attainment and the optimality of the paths taken. Among methods adhering to the same open-vocabulary, mapless, and universal constraints, our ImagineNav++ delivers markedly better performance than PSL [64], more than doubling both SR (52.4% vs. 23.0%) and SPL (32.8% vs. 11.4%). These results highlight the capability of our ImagineNav++ to enable efficient and robust navigation relying solely on visual observations, without requiring explicit map construction. The efficacy of our framework on instance‑level objectives further validates the strong adaptability of our ImagineNav++ across diverse goal‑oriented visual navigation tasks.
@@ IV-F1 Choice of Imagination Horizon
Figure 5: Where2Imagine: the impact of different sampling intervals $\mathbf{T}$ on ObjectNav performance on HM3D dataset.
@@
We evaluate the impact of different visual backbones on both waypoint prediction and final navigation performance. All backbones were adapted for our task: ResNet-18 and ViT were modified at their final layers and trained end-to-end, whereas DINOv2 and MAE were kept frozen, with their features fed into a separate trainable five-layer MLP. All experiments used real RGB observations without resorting to the NVS model and were tested for 100 epochs. As shown in Table IV, ResNet-18 achieves the best performance in both waypoint prediction and ObjectNav, despite being the most lightweight architecture. Furthermore, the significant performance gap across different backbones underscores the importance of our Where2Imagine module, indicating the usefulness of learning from human demonstrations. We also note a slight discrepancy between the waypoint prediction loss and the navigation success rate for DINOv2 and MAE. This is likely attributable to the domain shift between the MP3D dataset (used for waypoint evaluation) and the HM3D environment (used for navigation), which introduces evaluation variability.
@@
Figure 7 presents a comprehensive qualitative evaluation of our proposed Where2Imagine module on the HM3D dataset. This visualization reveals several key insights. First, the waypoints generated by our Where2Imagine module are concentrated in navigable regions, yielding observation views that are semantically richer and capture essential structural elements such as furniture, doors, and hallways. This stands in clear contrast to the uniform sampling strategy (with a fixed radius of 2.0m and angular interval of $60^{\circ}$), which disregards scene geometry and frequently fails to capture critical semantic information. This observation confirms the capability of our Where2Imagine module for semantic-aware planning, aligning its exploratory behavior with perceptually salient structural cues. Another key finding emerges from comparing the ground-truth and synthesized view observations. Specifically, although the NVS model is prone to introducing artifacts and inaccuracies, notably in the synthesis of fine details, they generally achieve fundamental consistency with the true scene observation in both semantics and overall structure across diverse HM3D environments. This high-level consistency enables effective spatial reasoning in LLMs, which explains the performance gains observed in navigation tasks.
@@ IV-H Analysis of VLM planner
We conducted a comparative evaluation of the effects of different VLMs on navigation performance, as detailed in Table VI. The experiments used real RGB without NVS model. The results demonstrate that GPT-4o-mini [50] and GPT-4-Turbo [49] exhibit negligible differences in success rate and SPL metrics, while substantially outperforming LLaVA [43], demonstrating that advanced reasoning capabilities are crucial for navigation tasks. Furthermore, our results indicate that within the same model family, strategically selecting more cost-effective variants (such as GPT-4o-mini vs. GPT-4-Turbo) achieves comparable navigation performance while dramatically reducing operational costs by approximately 98%. This optimization enables more efficient resource allocation and faster inference without sacrificing navigation performance, providing practical guidance for real-world deployment scenarios where computational efficiency and operational costs are significant concerns.
@@
GPT-4-Turbo [49] \$20.00 66.0 36.3
GPT-4o-mini [50] \$0.375 70.0 36.9
@@ IV-J Analysis of Successful and Failed Trajectories
As shown in Figure 10, our method enables efficient path planning and navigation toward diverse targets. In particular, the top-middle subfigure illustrates a scenario where the agent must traverse multiple rooms. Although such complex environments increase the risk of disorientation, our ImagineNav++ successfully reaches the goal after exploration, demonstrating its robustness in long-horizon, multi-room settings. Furthermore, as depicted in the upper-right subfigure, the memory module helps detect local search loops from historical observations and autonomously adjust the exploration direction. We also present some failure examples at the bottom of Figure 10. We identified three key factors contributing to these navigation failures. First, the synthesized image from the NVS does not align with the real observation, such as creating objects that are not present in the real scene as shown in the bottom left of Figure 10, which causes the VLM to make incorrect inference. Second, some object instances are neglected for marking by the simulator, and therefore a successful trajectory is wrongly considered as a failure (a.k.a. false failure) as shown in the middle of Figure 10. Moreover, the visual ambiguity of the target leads directly to recognition failure, causing the agent to terminate the navigation task unsuccessfully due to the lack of a key basis for the stopping decision, as shown in the bottom right of Figure 10.